• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 16:33:22 on August23,2003

    Entry number 112942

    keyword=day shift summary + bta

    day shift run summary

    Beam was great! Every now and then the pos feedback would turn itself off. The position however was wonderful all shift. We had another couple of auto start run fails for the right arm. We also discovered that for some kinematics, we weren't correcting for background dilution properly.


    0800:0901 kin 12.12, pol 3He, 270 deg / IN (still going)
    0816:0853 MCC called: dropped RF zone
    0901:0926 rotate to 0 deg, NMR = 31.9%
    0926:1137 kin 12.12, pol 3He, 0 deg / IN
    1137:1139 BHWP = OUT
    1139:1349 kin 12.12, pol 3He, 0 deg / OUT
    1349:1430 changing to kin 12.13, move to ref cell, NMR = 32.7% + EPR = 32.3%
    1430:1438 kin 12.13, ref cell = empty / OUT
    1438:1443 filling with N2 w/ 141 psig
    1443:1451 kin 12.13, ref cell = N2 / OUT
    1451:1501 move to Pol 3He
    1501:1600 kin 12.13, pol 3He, 0 deg / OUT



    FIGURE 1